OpenStudy (anonymous):

Finding Equilibrium pressure and final mass of a solid in a gas equation, comprehensive help?

OpenStudy (anonymous):

In the following equation: \[I_{2}(l) + H_{2}(g) \leftarrow \rightarrow 2HI(g)\] Given: Temperature = 25 celcius = 25+273 = 298 Kelvins Kc = 200

OpenStudy (anonymous):

So I found this later: Kp = 4893 The starting pressure = 5.00 atm Volume = 2.00 L Starting moles of I2 = twice of H2 (I found the mass of I2 starting to be 208g) SO my problem is, how do I find the 1. Equilibrium Pressures and more pressingly 2. Final mass of I2 at equillibrium.

OpenStudy (anonymous):

Here's my take. Since I already found Kp, it is equal to P of products / P of Reactants. since only 1 on both sides... I already found the initial P of it, which is 5.00. The change of the ICE box is so... 4893 = 4x^2/5-x Getting the quadratic \[4x^{2} + 4893x - 24463\] Which I found x = 4.97 Final pressure: 5 -4.97 = 0.03 H2 2x4.97 = 9.94 HI